NASA ISS On-Orbit Status 19 September 2013

©NASA

NASA ISS On-Orbit Status 19 September 2013.

Cygnus Status: Cygnus completed its far-field demonstrations today. Delivery of the data package for Demo 1 (Position and Attitude Control) to NASA has been delayed due to TDRS communications issues.

However, is expected to be delivered to NASA this evening. Data packages for Demo 2a & 2b (Free Drift & Active Abort) were delivered to NASA this morning. NASA teams have reviewed the information and determined that all objectives for Demos 2a & 2b have been satisfied. NASA teams will review the Demo 1 information this evening, prior to tomorrow's IMMT Go/No-Go poll for Joint Operations. The DV3 phasing burn was rescheduled to 4:22 pm CDT today due to a post-burn conjunction concern. TDRS drop-outs and ratty comm continue, but Cygnus systems remain nominal. Capture and berthing to the Node 2 Nadir port is planned for Sunday, 22 September.

Ocular Health: FE-5 Parmitano and FE-6 Nyberg completed their three day series of Ocular Health activities for their flight day 120 session. This was their fifth of six planned sessions. Alternating as a Crew Medical Officer (CMO) and as a subject, Parmitano and Nyberg performed ocular and cardiac ultrasound scans and blood pressure measurements. In-flight data collection was assisted by remote guidance from the Ground to ensure proper positioning and data collection. Ocular ultrasound will be used to identify changes in globe morphology, including flattening of the posterior globe, and document optic nerve sheath diameter, optic nerve sheath tortuosity, globe axial measurements, and choroidal engorgement. The Ocular Health protocol calls for a systematic gathering of physiological data to characterize the risk of microgravity-induced visual impairment/intracranial pressure in ISS crewmembers. Researchers believe that the measurement of visual, vascular and central nervous system changes over the course of this experiment and during the subsequent postflight recovery will assist in the development of countermeasures, clinical monitoring strategies and clinical practice guidelines.

Circadian Rhythms: Today Parmitano secured Thermolab sensors to himself and attached the Thermolab Unit to his belt. The Equipment will be worn for the next 36 hours. This is the fifth of six sessions planned for Parmitano. The ESA Circadian Rhythms experiment examines the role of synchronized circadian rhythms (the human body's 24 hour light-dark cycle) and its maintenance during long duration spaceflight. Understanding how changes in body composition, body temperature regulation and physical activity affect circadian rhythms will support development of enhancements for performance and health care of future crewmembers.

ISS SERVIR Environmental Research and Visualization System (ISERV) Troubleshooting: As part of the troubleshooting for ISERV, Nyberg accessed the Window Obervational Research Facility (WORF) and performed a telescope motion configuration with the ISERV hand controller. Upon ground commanding from the ISERV operations center, she observed that the pointing motion was not nominal. She then manually pointed the telescope in a nadir direction for near term observations. Her observations confirmed that the pointing anomaly is not software related. Instead it is due to a failure in the pointing system's encoder. Teams will continue their nadir imaging operations. ISERV is expected to provide images for use in disaster monitoring and assessment, and environmental decision making.

Today's Planned Tasks

PROK- Urine Sampling [Complete]
HRF - urine sampling [Complete]
HRF - sample insertion into MELFI [Complete]
USND2- hw test before experiment [Complete]
URAGAN. Charging ФСС batteries (start) [Complete]
Comm config for MRM2 experiment support [Deferred]
EPM - laptop activation [Complete]
KULONOVSKIY KRISTAL. Experiment Ops. [Complete]
ECLSS Recycle Tank Remove and Replace [Complete]
HRF - urine sampling [Complete]
HRF - insterting samples into MELFI [Complete]
Ultrasound- scan prep [Complete]
Ocular Health (OH) Ultrasound 2 Scan [Complete]
Comm reconfig in MRM2 [Deferred]
KULONOVSKIY KRISTAL. Data copy to hard drive [Complete]
Ultrasound - data downlink [Complete]
Ocular Health (OH) Ultrasound 2 Scan [Complete]
SHD - weekly questionnaire [Complete]
Ultrasound - post-test activities [Complete]
Ultrasound - data downlink [Complete]
OH - heart scan and blood pressure reading [Complete]
СОЖ maintenance [Complete]
Recycle Tank Drain into ATV [Complete]
IDENTIFIKATSIA ИМУ-Ц results data copying to laptop [Complete]
ISERV - ISS SERVIR - Evironment study and visualization system [Complete]
ATV - air repress start [Complete]
Robotic Workstation (RWS) Setup [Complete]
OH - heart scan and blood pressure reading [Complete]
АСУ - R&Rs [Complete]
Robotic Workstation (RWS) High Definition (HD) Monitor Downlink [Complete]
USND2- final steps and hw stow [Complete]
PROK- prep and diet menu recording [Complete]
IMS and stowage conference [Complete]
HRF - urine sampling [Complete]
HRF - sample insertion into MELFI [Complete]
O2/CO2 sampling [Complete]
URAGAN Observation and photography using ФСС equipment [Complete]
ATV - air repress (end) [Complete]
HRF - urine sampling [Complete]
HRF - sample insertion into MELFI [Complete]
HRF- hw install [Complete]
Robotic Workstation (RWS) Monitor HD Video Downlink Deactivation [Complete]
Thermolab Instrumentation for Circadian Rhythms [Complete]
USOS window shutters close [Complete]
Blood sampling reminder [Complete]
Task List

ARED Cylinder Evacuation
IMS edit
КТО replace
ЕДВ-У R&R
Preparation of reports for Roskosmos site
URAGAN. Observations and Photography
ECON-M. Observations and Photography
WSTA Fill
FFQ [Complete]
ACS Cable Install [Complete]
DIAPASON CAROUSEL #3 RUN [Complete]
Install Covers over the Crew Quarters Blankets [In Work]
T2 Snubber Filing [Complete]
Ground

Nominal Ground Commanding [Complete]
Service Module (SM) Solar Array Efficiency Test
Three-Day Look Ahead:
Friday, 9/20: HRF Blood/Urine collection, Cygnus Vehicle Ops OBT
Saturday, 9/21: SSRMS DOUG Setup, Off Duty
Sunday, 9/22: Cygnus Rendezvous, Capture, and Install

QUICK ISS Status - Environmental Control Group:

Component - Status
Elektron - On
Vozdukh - Auto
[СКВ] 1 - SM Air Conditioner System ("SKV1") - Off
[СКВ] 2 - SM Air Conditioner System ("SKV2") - Off
Carbon Dioxide Removal Assembly (CDRA) Lab - Standby
Carbon Dioxide Removal Assembly (CDRA) Node 3 - Operate
Major Constituent Analyzer (MCA) Lab - Off
Major Constituent Analyzer (MCA) Node 3 - Operate
Oxygen Generation Assembly (OGA) - Process
Urine Processing Assembly (UPA) - Standby
Trace Contaminant Control System (TCCS) Lab - Off
Trace Contaminant Control System (TCCS) Node 3 - Full Up


Please follow SpaceRef on Twitter and Like us on Facebook.